Gaseous ethane CH3CH3 will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O. Suppose 0.601 g of ethane is mixed with 1.7 g of oxygen. Calculate the maximum mass of carbon dioxide that could be produced by the chemical reaction. Round your answer to 2 significant digits.


1
Expert's answer
2021-03-08T06:05:44-0500

2CH3CH3 + 7O2 ----> 4CO2 + 6H2O

Moles of ethane = 0.601/30

= 0.020 mol

Moles of O2 = 1.7/32

= 0.053 mol

O2 is limiting reagent.

Mass of CO2 produced = 4×44×1.7/7×32

= 1.3 g

Mass of carbondioxide = [1.3] g